A changing current induces an E.M.F of 10v across a 0.25H inductor. What is the rate at which the current is changing?


1
Expert's answer
2021-08-30T15:05:25-0400

 We define the self-inductance of the circuit as "L=\frac{N\Phi_B}{i}" and the units for the inductance are "1\,H =1\frac{V\cdot s}{A}".


From Faraday’s law for a coil with N turns, the self-induced emf is "\epsilon = - N \frac{d \Phi_B}{dt}". Using the prior relation we have:


"\epsilon = - N \frac{d \Phi_B}{dt}=-L\frac{d i}{dt} \implies |\epsilon|=L|\frac{d i}{dt}|
\\ |\frac{d i}{dt}|=\cfrac{|\epsilon|}{L}=\cfrac{|10\,V|}{0.25\,H}=\cfrac{10\,V}{0.25\frac{V\cdot s}{A}}
\\ |\frac{d i}{dt}|=40 \frac{A}{s}"


In conclusion, the rate of change for the changing current is 40 A/s.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.